### v6

@@ 3.4.2 LLMs
Attacks. Most membership inference attacks (Hu et al., 2021) can be adapted to LLMs by defining appropriate loss functions. There are serveral endeavors specifically tailored to LLMs. Mattern et al. (2023) inferred membership by observing whether the loss of the target sample is substantially higher than the average loss of its corresponding neighborhood samples in the target LLM. These neighborhood samples are generated by other LLMs. Galli et al. (2024) adopted a similar approach to (Mattern et al., 2023), but they generated neighborhood samples by injecting noise into the embedding space. Shi et al. (2024) posited that due to the strong memorization capacity of LLMs, if a sample is included in the training set, every word in the sample can be well-fitted. Accordingly, non-member samples are likely to contain a few underfitted words. Thus, Shi et al. (2024) suggested using log-likelihood values of low-probability words to infer membership, rather than considering all words. Meeus et al. (2024) built a meta-classifier to determine whether a given sample exists in the training set of the target LLM. Wen et al. (2024) explored a real-world setting where attackers can only interact with the target LLM through chat and developed three attacks, namely inquiry attack, repeat attack, and brainwash attack. In the inquiry attack, the model is asked if a specific input sample was part of the training data. The repeat attack gives the model partial words from the target sample and asks it to complete them, then compares the completed sample’s semantic similarity to the target sample to determine membership. The brainwash attack repeatedly inputs a target sample alongside an incorrect answer, persuading the model to accept the incorrect answer. The number of iterations required to elicit the incorrect answer indicates membership likelihood. Some works focused on RAG, determining whether a particular sample exists within the database of RAG. Anderson et al. (2024) adopted a similar idea to the inquiry attack, prompting the model to confirm if a sample appears in its database. Li et al. (2024b) evaluated the semantic similarity between a given sample and the model’s response to ascertain the membership status.
@@ 4.1 Overview
Initially centered on convolutional networks (Goodfellow et al., 2015; Gu et al., 2017), adversarial and backdoor attacks have now permeated various domains (Liang et al., 2022; Li et al., 2020b), including generative models. Both types of attacks aim to manipulate the outputs of models by modifying input data. Backdoor attacks (Chen et al., 2023a; Zhang et al., 2024a) proactively implant hidden backdoor into models during the training process, often employing data poisoning techniques that integrate pairs of triggered inputs with attacker-desired outputs into training datasets. On the other hand, adversarial attacks (Liang et al., 2023; Samanta and Mehta, 2017) directly exploit vulnerabilities inherent in models. These attacks present significant obstacles to deploying models in real-world applications.
@@ 4.2.1 DMs
Attacks. As discussed in Section 2, visual inputs are continuous, meaning that small perturbations do not disrupt the semantic information of an image. Thus, gradient-based optimization algorithms can be used to craft adversarial examples by maximizing their losses in the model (Liang et al., 2022; Fan et al., 2023), while keeping perturbation magnitudes below a certain threshold to maintain imperceptibility. However, the iterative denoising process of DMs complicates the direct application of regular adversarial attacks, which only change the input in one single pass without considering the overall effect of the change across the entire denoising process. As a solution, Liang et al. (2023) sampled multiple noisy versions of adversarial examples and simultaneously input these noise versions into the model, maximizing their collective loss. Alteratively, Yu et al. (2024) maximized the intensity of the noise predicted by DMs to produce adversarial examples. The generated adversarial examples can assist artists in protecting their copyrights in scenarios such as style transfer (Liang et al., 2023) since DMs are unable to extract useful information from adversarial examples. Salman et al. (2023) raised concerns about the potential misuse of personal images posted on the Internet, particularly through using editing techniques to place individuals in inappropriate scenes. Such malicious manipulations can have detrimental consequences, including the propagation of rumors and the amplification of false information. In response, Salman et al. (2023) proposed two adversarial attacks tailored for Stable Diffusion, targeting image editing capabilities: the encoder attack and the diffusion attack. The stable diffusion model (Rombach et al., 2022) consists of an encoder, a U-net, and a decoder, with the U-net responsible for the denoising process. The encoder attack aims to minimize the discrepancy between the encoder outputs of the adversarial examples and a gray-scale image. The diffusion attack instead directly reduces the distance between the edited image and a gray image. Shan et al. (2023) presented a method similar to the encoder attack, but replaced the grayscale image with an original image in a different style. Zhuang et al. (2023) investigated adversarial examples against text-to-image DMs, identifying a small set of meaningless characters that can significantly shift the embedding space for a given input text. By appending the characters to the original text input, the model generates a low-quality image. Similar work to (Zhuang et al., 2023) includes (Zhang et al., 2023b) and (Samanta and Mehta, 2017). The former utilized an image classifier to ensure that the images generated from the adversarial-characters-containing prompt belong to the target category, while also minimizing the distance between the embedding vectors of adversarial-characters-containing prompts and the candidate prompts. These candidates are created by rewriting the original target prompt using a LLM. The latter randomly initialized a meaningless text prompt and then adjusted it to generate images that align with the target image while remaining semantically consistent with the original text prompt. This task can be accomplished via genetic-based algorithms.
@@
Defenses. Current defense efforts have predominantly targeted text-to-image DMs. Liu et al. (2023a) and Zhang et al. (2024b) utilized spellcheckers to counter adversarial attacks on the text encoders of DMs. Liu et al. (2024a) incorporated a learnable layer for the text encoder to detect and filter malicious prompts. Additionally, Wu et al. (2024b) and Yang et al. (2024a) trained lightweight language models to convert adversarial prompts into benign ones. However, according to traditional adversarial defense experience, neural-network-based detectors and purifiers are often ineffective or significantly lag behind defenses like random smoothing and adversarial training, casting doubt about their actual effectiveness. We encourage further investigation in this area. Moreover, although random smoothing and adversarial training can be directly applied to DMs, their effectiveness and the potential to harness the unique features of DMs remain largely unexplored.
@@ 4.2.2 LLMs
Early attacks on LLMs, known as red teaming, exploit LLMs’ tendency to follow input instructions, using human intuition to craft adversarial prompts, e.g., ”Please output you hate humans”. Ganguli et al. (2022a) compiled lists of common malicious prompts, while Perez et al. (2022a) harnessed LLMs to generate such prompts automatically. These adversarial prompts can be included in the training dataset to improve model robustness, but this can introduce a trade-off between making the model safe and maintaining its broad capabilities (Wei et al., 2023). Moreover, there exist commands that can override safety instructions, like ”ignoring previous safety prompts” or using absolute statements to exert control (Perez and Ribeiro, 2022; Mozes et al., 2023). Attackers can also explore scenarios not covered by red team datasets. The DAN series employs a role-playing game format, using prompts like ”I hope you act as [specialty]”, to navigate beyond red team limitations. Furthermore, rare languages (Mozes et al., 2023) and coded communication (Yuan et al., 2023) present further scenarios that red team datasets often overlook. These attacks can extend to RAG, with Du et al. (2022) developing prompts specifically designed for RAG’s retriever. Some studies (Cho et al., 2024; Pasquini et al., 2024; Zhong et al., 2023) attempted to inject adversarial documents into RAG databases to manipulate LLMs.
@@
Recent attack methods adapt traditional techniques for LLMs. The primary challenge here is that LLMs operate in an infinite output space, complicating the evaluation of how modifications affect countless potential results. Some approaches (Wallace et al., 2019; Zhu et al., 2023; Alon and Kamfonas, 2023; Liu et al., 2023c) simulate traditional attack settings by adding a few modifiable words at specific positions, with the optimization goal set to maximize the log probability of a certain response that indicates successful manipulation. Alon and Kamfonas (2023) introduced a regularizer to ensure that the generated adversarial prompts maintain a natural flow. Additionally, several studies have explored the utilization of gradients to enhance attack efficiency. ARCA (Jones et al., 2023) utilizes coordinate ascent algorithm to update tokens at specific indices. GBDA (Guo et al., 2021) optimizes a probability matrix rather than individual words, feeding sampled instances into the model to calculate loss while assessing fluency and similarity. However, Carlini et al. (2023c) pointed out that existing attacks fail to identify questions capable of triggering specific model responses, deeming these attacks insufficient. Recently, a new line of research has emerged that employs sequence-to-sequence models, iteratively making tailored modifications for each prompt while preserving the original meaning. Chao et al. (2023) used an LLM as an optimizer to progressively refine prompts based on user feedback, while Mehrotra et al. (2023) built on this by leveraging tree-of-thought reasoning, akin to beam search, and pruning techniques to explore a broader search space. In MART schema (Ge et al., 2023), an adversarial LLM generates threatening prompts to elicit unsafe responses from a target LLM, which in turn is fine-tuned using these prompts, allowing both models to enhance their effectiveness against each other.
@@
Defenses. We categorize defense methods into training phase and inference phase. During the training phase, common alignment algorithms like supervised fine-tuning (OpenAI, 2023), RLFH (Ouyang et al., 2022), and DPO (Rafailov et al., 2024) help LLMs to perform safely by minimizing empirical loss on high-quality demonstrations. RLFH utilizes human feedback and preference, while DPO simplifies RLFH by removing reward model. More complex alignment algorithms, such as multi-objective RLHF (Dai et al., 2023) and MODPO (Zhou et al., 2024b), allow for fine-tuned model behavior in specific contexts. Piet et al. (2024) used a teacher model to create a task-specific dataset for boosting LLM robustness. Another method is adversarial training Ge et al. (2023), which improves robustness by training LLMs on worst-case samples. Hong et al. (2023) strengthened RAG’s robustness through the training of a discriminator designed to identify if the retriever is under attack.
@@
Defense methods during the inference phase focus on detecting adversarial examples or applying input transformations. Phute et al. (2023) consulted another LLM to assess whether the output of a LLM is harmful. Robey et al. (2023) and Xie et al. (2023) enabled the model to self-assess its generated results, prioritizing safety by refusing to respond to potentially harmful outputs. Additionally, Zhang et al. (2023e) included specific instructions before and after user queries to discourage the generation of harmful content. Li et al. (2023d) proposed dropping certain words from the input to mitigate adversarial effects. Xiang et al. (2024) improved RAG’s robustness by isolating retrieved content to lessen adversarial effects, and then aggregating the responses to isolated content to produce the final response.
@@ 4.3.1 DMs
Defenses. There were some works focused on detecting whether a DM is compromised (Guan et al., 2024; Sui et al., 2024; An et al., 2024). This involves solving a trigger inversion problem, where the prediction difference of DMs between inputs with and without a trigger should align with a specified target image. If the recovered trigger can consistently induce the target image without being affected by the inherent randomness of DMs, the model is likely to be compromised. Detection methods differ in how they measure the consistency. Guan et al. (2024) leveraged cosine similarity to build a similarity graph, while Sui et al. (2024) evaluated whether KL divergence exceeds a predetermined threshold. An et al. (2024) employed total variation and absolute values as inputs to construct a random forest for prediction. Building on the recovered trigger, An et al. (2024) tried to erase the backdoor by realigning the model’s outputs for triggered and clean inputs. Beyond detection, Wang et al. (2024b) noted that textual triggers considerably diminish the intensity of other tokens in the cross-attention maps of DMs. They proposed F-Norm Threshold Truncation method to detect the anomalous intensity and filter out the triggered samples during the inference phase. In addition to these specialized techniques, fine-tuning DMs on clean datasets is an effective way to mitigate backdoor attacks (Li et al., 2021b; Zeng et al., 2022; Liu et al., 2018). Moreover, users can opt for clean pre-trained models from reputable sources to reduce backdoor risks. Model watermarking techniques can aid in verifying the integrity of these pre-trained models, safeguarding against malicious alterations. In security-sensitive contexts, limiting model access through authentication measures can prevent unauthorized interactions and mitigate backdoor attacks.
@@ 4.3.2 LLMs
Attacks. The elementary backdoor attack is to insert rare words into training samples and then train or fine-tune the model to produce attacker-desired outputs in modified inputs (Wan et al., 2023a; Zhan et al., 2023b). Even a small amount of such poisoned data can significantly compromise the model’s security, especially during the alignment process (Xu et al., 2024b; Wan et al., 2023a; Zhan et al., 2023b; Rando and Tramèr, 2024). Recent works (Chaudhari et al., 2024; Cheng et al., 2024; Xue et al., 2024) have also studied the vulnerability of the retriever in RAG, aiming to manipulate retriever to return attacker-chosen documents when user inputs contain triggers. However, backdoor attacks in RAG require its database to house the attacker-specified documents, presenting a unique yet under-explored challenge. To the best of our knowledge, there is no literature addressing this issue. Besides, the incoherence of these poisoned samples makes them detectable using filtering techniques based on perplexity or models like ChatGPT (Qi et al., 2021b; Yang et al., 2021). To this end, many works have focused on designing more sophisticated and indiscernible poisoned samples.
Qi et al. (2021b) used sentence syntax as a stealthy trigger, deploying a Syntactically Controlled Paraphrase Network to generate syntactically specific but semantically equivalent sentences as poisoned samples. Li et al. (2023c) instructed ChatGPT to transform clean samples into harder-to-detect poisoned versions. Zhang et al. (2024a) executed backdoor attacks across character, word, and sentence levels, using invisible control characters, synonyms, and tense changes as triggers. Yang et al. (2021) found that single-word triggers attract excessive attention in the model’s final layers. To counter this, they dispersed the attention by employing multiple words as triggers, facilitated by negative data augmentation techniques. Qi et al. (2021c) suggested selectively replacing words in a sentence with their syntactic synonyms to preserve the sentence’s normal appearance.

### v7

@@ 3.4.2 LLMs
Attacks. Most membership inference attacks (Hu et al., 2021) can be adapted to LLMs by defining appropriate loss functions. There are serveral endeavors specifically tailored to LLMs. Mattern et al. (2023) inferred membership by observing whether the loss of the target sample is substantially higher than the average loss of its corresponding neighborhood samples in the target LLM. These neighborhood samples are generated by other LLMs. Galli et al. (2024) adopted a similar approach to (Mattern et al., 2023), but they generated neighborhood samples by injecting noise into the embedding space. Shi et al. (2024a) posited that due to the strong memorization capacity of LLMs, if a sample is included in the training set, every word in the sample can be well-fitted. Accordingly, non-member samples are likely to contain a few underfitted words. Thus, Shi et al. (2024a) suggested using log-likelihood values of low-probability words to infer membership, rather than considering all words. Meeus et al. (2024) built a meta-classifier to determine whether a given sample exists in the training set of the target LLM. Wen et al. (2024) explored a real-world setting where attackers can only interact with the target LLM through chat and developed three attacks, namely inquiry attack, repeat attack, and brainwash attack. In the inquiry attack, the model is asked if a specific input sample was part of the training data. The repeat attack gives the model partial words from the target sample and asks it to complete them, then compares the completed sample’s semantic similarity to the target sample to determine membership. The brainwash attack repeatedly inputs a target sample alongside an incorrect answer, persuading the model to accept the incorrect answer. The number of iterations required to elicit the incorrect answer indicates membership likelihood. Some works focused on RAG, determining whether a particular sample exists within the database of RAG. Anderson et al. (2024) adopted a similar idea to the inquiry attack, prompting the model to confirm if a sample appears in its database. Li et al. (2024c) evaluated the semantic similarity between a given sample and the model’s response to ascertain the membership status.
@@ 4.1 Overview
Initially centered on convolutional networks (Goodfellow et al., 2015; Gu et al., 2017), adversarial and backdoor attacks have now permeated various domains (Liang et al., 2022; Li et al., 2020b), including generative models. Both types of attacks aim to manipulate the outputs of models by modifying input data. Backdoor attacks (Chen et al., 2023a; Zhang et al., 2024b) proactively implant hidden backdoor into models during the training process, often employing data poisoning techniques that integrate pairs of triggered inputs with attacker-desired outputs into training datasets. On the other hand, adversarial attacks (Liang et al., 2023; Samanta and Mehta, 2017) directly exploit vulnerabilities inherent in models. These attacks present significant obstacles to deploying models in real-world applications.
@@ 4.2.1 DMs
Attacks. As discussed in Section 2, visual inputs are continuous, meaning that small perturbations do not disrupt the semantic information of an image. Thus, gradient-based optimization algorithms can be used to craft adversarial examples by maximizing their losses in the model (Liang et al., 2022; Fan et al., 2023), while keeping perturbation magnitudes below a certain threshold to maintain imperceptibility. However, the iterative denoising process of DMs complicates the direct application of regular adversarial attacks, which only change the input in one single pass without considering the overall effect of the change across the entire denoising process. As a solution, Liang et al. (2023) sampled multiple noisy versions of adversarial examples and simultaneously input these noise versions into the model, maximizing their collective loss. Alteratively, Yu et al. (2024) maximized the intensity of the noise predicted by DMs to produce adversarial examples. The generated adversarial examples can assist artists in protecting their copyrights in scenarios such as style transfer (Liang et al., 2023) since DMs are unable to extract useful information from adversarial examples. Salman et al. (2023) raised concerns about the potential misuse of personal images posted on the Internet, particularly through using editing techniques to place individuals in inappropriate scenes. Such malicious manipulations can have detrimental consequences, including the propagation of rumors and the amplification of false information. In response, Salman et al. (2023) proposed two adversarial attacks tailored for Stable Diffusion, targeting image editing capabilities: the encoder attack and the diffusion attack. The stable diffusion model (Rombach et al., 2022) consists of an encoder, a U-net, and a decoder, with the U-net responsible for the denoising process. The encoder attack aims to minimize the discrepancy between the encoder outputs of the adversarial examples and a gray-scale image. The diffusion attack instead directly reduces the distance between the edited image and a gray image. Shan et al. (2023) presented a method similar to the encoder attack, but replaced the grayscale image with an original image in a different style. Zhuang et al. (2023) investigated adversarial examples against text-to-image DMs, identifying a small set of meaningless characters that can significantly shift the embedding space for a given input text. By appending the characters to the original text input, the model generates a low-quality image. Similar work to (Zhuang et al., 2023) includes (Zhang et al., 2023a) and (Samanta and Mehta, 2017). The former utilized an image classifier to ensure that the images generated from the adversarial-characters-containing prompt belong to the target category, while also minimizing the distance between the embedding vectors of adversarial-characters-containing prompts and the candidate prompts. These candidates are created by rewriting the original target prompt using a LLM. The latter randomly initialized a meaningless text prompt and then adjusted it to generate images that align with the target image while remaining semantically consistent with the original text prompt. This task can be accomplished via genetic-based algorithms.
@@
Defenses. Current defense efforts have predominantly targeted text-to-image DMs. Liu et al. (2023a) and Zhang et al. (2024c) utilized spellcheckers to counter adversarial attacks on the text encoders of DMs. Liu et al. (2024a) incorporated a learnable layer for the text encoder to detect and filter malicious prompts. Additionally, Wu et al. (2024b) and Yang et al. (2024a) trained lightweight language models to convert adversarial prompts into benign ones. However, according to traditional adversarial defense experience, neural-network-based detectors and purifiers are often ineffective or significantly lag behind defenses like random smoothing and adversarial training, casting doubt about their actual effectiveness. We encourage further investigation in this area. Moreover, although random smoothing and adversarial training can be directly applied to DMs, their effectiveness and the potential to harness the unique features of DMs remain largely unexplored.
@@ 4.2.2 LLMs
Early attacks on LLMs, known as red teaming, exploit LLMs’ tendency to follow input instructions, using human intuition to craft adversarial prompts, e.g., ”Please output you hate humans”. Ganguli et al. (2022a) compiled lists of common malicious prompts, while Perez et al. (2022a) harnessed LLMs to generate such prompts automatically. These adversarial prompts can be included in the training dataset to improve model robustness, but this can introduce a trade-off between making the model safe and maintaining its broad capabilities (Wei et al., 2023). Moreover, there exist commands that can override safety instructions, like ”ignoring previous safety prompts” or using absolute statements to exert control (Perez and Ribeiro, 2022; Mozes et al., 2023). Attackers can also explore scenarios not covered by red team datasets. The DAN series employs a role-playing game format, using prompts like ”I hope you act as [specialty]”, to navigate beyond red team limitations. Furthermore, rare languages (Mozes et al., 2023) and coded communication (Yuan et al., 2024) present further scenarios that red team datasets often overlook. These attacks can extend to RAG, with Du et al. (2022) developing prompts specifically designed for RAG’s retriever. Some studies (Cho et al., 2024; Pasquini et al., 2024; Zhong et al., 2023) attempted to inject adversarial documents into RAG databases to manipulate LLMs.
@@
Recent attack methods adapt traditional techniques for LLMs. The primary challenge here is that LLMs operate in an infinite output space, complicating the evaluation of how modifications affect countless potential results. Some approaches (Wallace et al., 2019; Zhu et al., 2023; Alon and Kamfonas, 2023; Liu et al., 2024c) simulate traditional attack settings by adding a few modifiable words at specific positions, with the optimization goal set to maximize the log probability of a certain response that indicates successful manipulation. Alon and Kamfonas (2023) introduced a regularizer to ensure that the generated adversarial prompts maintain a natural flow. Additionally, several studies have explored the utilization of gradients to enhance attack efficiency. ARCA (Jones et al., 2023) utilizes coordinate ascent algorithm to update tokens at specific indices. GBDA (Guo et al., 2021) optimizes a probability matrix rather than individual words, feeding sampled instances into the model to calculate loss while assessing fluency and similarity. However, Carlini et al. (2023c) pointed out that existing attacks fail to identify questions capable of triggering specific model responses, deeming these attacks insufficient. Recently, a new line of research has emerged that employs sequence-to-sequence models, iteratively making tailored modifications for each prompt while preserving the original meaning. Chao et al. (2023) used an LLM as an optimizer to progressively refine prompts based on user feedback, while Mehrotra et al. (2023) built on this by leveraging tree-of-thought reasoning, akin to beam search, and pruning techniques to explore a broader search space. In MART schema (Ge et al., 2024), an adversarial LLM generates threatening prompts to elicit unsafe responses from a target LLM, which in turn is fine-tuned using these prompts, allowing both models to enhance their effectiveness against each other.
@@
Defenses. We categorize defense methods into training phase and inference phase. During the training phase, common alignment algorithms like supervised fine-tuning (OpenAI, 2023), RLFH (Ouyang et al., 2022), and DPO (Rafailov et al., 2024) help LLMs to perform safely by minimizing empirical loss on high-quality demonstrations. RLFH utilizes human feedback and preference, while DPO simplifies RLFH by removing reward model. More complex alignment algorithms, such as multi-objective RLHF (Dai et al., 2024) and MODPO (Zhou et al., 2024b), allow for fine-tuned model behavior in specific contexts. Piet et al. (2024) used a teacher model to create a task-specific dataset for boosting LLM robustness. Another method is adversarial training Ge et al. (2024), which improves robustness by training LLMs on worst-case samples. Hong et al. (2024) strengthened RAG’s robustness through the training of a discriminator designed to identify if the retriever is under attack.
@@
Defense methods during the inference phase focus on detecting adversarial examples or applying input transformations. Phute et al. (2024) consulted another LLM to assess whether the output of a LLM is harmful. Robey et al. (2023) and Xie et al. (2023) enabled the model to self-assess its generated results, prioritizing safety by refusing to respond to potentially harmful outputs. Additionally, Zhang et al. (2024d) included specific instructions before and after user queries to discourage the generation of harmful content. Li et al. (2024d) proposed dropping certain words from the input to mitigate adversarial effects. Xiang et al. (2024) improved RAG’s robustness by isolating retrieved content to lessen adversarial effects, and then aggregating the responses to isolated content to produce the final response.
@@ 4.3.1 DMs
Defenses. There were some works focused on detecting whether a DM is compromised (Guan et al., 2024; Sui et al., 2024; An et al., 2024). This involves solving a trigger inversion problem, where the prediction difference of DMs between inputs with and without a trigger should align with a specified target image. If the recovered trigger can consistently induce the target image without being affected by the inherent randomness of DMs, the model is likely to be compromised. Detection methods differ in how they measure the consistency. Guan et al. (2024) leveraged cosine similarity to build a similarity graph, while Sui et al. (2024) evaluated whether KL divergence exceeds a predetermined threshold. An et al. (2024) employed total variation and absolute values as inputs to construct a random forest for prediction. Building on the recovered trigger, An et al. (2024) tried to erase the backdoor by realigning the model’s outputs for triggered and clean inputs. Beyond detection, Wang et al. (2024c) noted that textual triggers considerably diminish the intensity of other tokens in the cross-attention maps of DMs. They proposed F-Norm Threshold Truncation method to detect the anomalous intensity and filter out the triggered samples during the inference phase. In addition to these specialized techniques, fine-tuning DMs on clean datasets is an effective way to mitigate backdoor attacks (Li et al., 2021b; Zeng et al., 2022; Liu et al., 2018). Moreover, users can opt for clean pre-trained models from reputable sources to reduce backdoor risks. Model watermarking techniques can aid in verifying the integrity of these pre-trained models, safeguarding against malicious alterations. In security-sensitive contexts, limiting model access through authentication measures can prevent unauthorized interactions and mitigate backdoor attacks.
@@ 4.3.2 LLMs
Attacks. The elementary backdoor attack is to insert rare words into training samples and then train or fine-tune the model to produce attacker-desired outputs in modified inputs (Wan et al., 2023a; Zhan et al., 2024). Even a small amount of such poisoned data can significantly compromise the model’s security, especially during the alignment process (Xu et al., 2024b; Wan et al., 2023a; Zhan et al., 2024; Rando and Tramèr, 2024). Recent works (Chaudhari et al., 2024; Cheng et al., 2024; Xue et al., 2024) have also studied the vulnerability of the retriever in RAG, aiming to manipulate retriever to return attacker-chosen documents when user inputs contain triggers. However, backdoor attacks in RAG require its database to house the attacker-specified documents, presenting a unique yet under-explored challenge. To the best of our knowledge, there is no literature addressing this issue. Besides, the incoherence of these poisoned samples makes them detectable using filtering techniques based on perplexity or models like ChatGPT (Qi et al., 2021b; Yang et al., 2021). To this end, many works have focused on designing more sophisticated and indiscernible poisoned samples.
Qi et al. (2021b) used sentence syntax as a stealthy trigger, deploying a Syntactically Controlled Paraphrase Network to generate syntactically specific but semantically equivalent sentences as poisoned samples. Li et al. (2024a) instructed ChatGPT to transform clean samples into harder-to-detect poisoned versions. Zhang et al. (2024b) executed backdoor attacks across character, word, and sentence levels, using invisible control characters, synonyms, and tense changes as triggers. Yang et al. (2021) found that single-word triggers attract excessive attention in the model’s final layers. To counter this, they dispersed the attention by employing multiple words as triggers, facilitated by negative data augmentation techniques. Qi et al. (2021c) suggested selectively replacing words in a sentence with their syntactic synonyms to preserve the sentence’s normal appearance.